. a locomotive of 60-tonne mass pulls some carriage of total mass 240 tonnes a distance of half a kilometers up a 1 in 200 inclines with an acceleration of 0.2m/s². if the overall rolling resistance is 5% of the normal reaction, what is the total work done?


1
Expert's answer
2022-02-17T10:05:33-0500

Given:

"m=\rm 60\: t+240\: t=300\: t=3*10^5\: kg"

"d=500\:\rm m"

"\sin\theta=1/200"

"a=0.2\:\rm m/s^2"

"\mu=0.05"


The net force


"F=ma=3*10^5*0.2=6*10^4\:\rm N"

The total work done


"W=Fd=6*10^4\:\rm N*500\: m=3*10^7\:\rm J"